OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–2, 1) and (4, 13)? What is the equation in point-slope form of the line passing through (–2, 1) and (4, 13)? @Mathematics

jimthompson5910 (jim_thompson5910):

slope: m = (y2-y1)/(x2-x1) m = (13-1)/(4-(-2)) m = (13-1)/(4+2) m = 12/6 m = 2 So the slope is m = 2 Now turn to the point slope formula y-y1 = m(x-x1) and plug in m = 2, x1 = -2 and y1 = 1 to get y-1 = 2(x-(-2)) y-1 = 2(x+2) which is your answer